# Universal neighborhood topology and geometry of exceptional points in physical systems

### v1

@@ Abstract
where $a=\delta+i\gamma+\alpha\left(|x|^{2}-|y|^{2}\right)$, $\alpha,\beta,\gamma,\delta\in\mathbb{R}$ are parameters, and all variables and parameters are assumed scaled and unitless. This parametric form is chosen for ease of analysis; the restrictions can be relaxed and the method is still applicable (see SM for details). In the linear limit, the two states are separated by $2\delta$ in frequency, and one state gains and the other dissipates at equal rates $|\gamma|$. The eigenvector, normalized to $1$, is written as $[x\,\,y]^{T}=(1/\textit{A})[1\,\,\tilde{w}]^{T}$, $\tilde{w}=we^{i\theta}$, $\textit{A}=\sqrt{1+w^{2}}$. Eliminating the eigenvalue $E$ from Eq. (1) reduces the matrix equation to a nonlinear equation in the complex eigenvector element $\tilde{w}$, the real and imaginary parts of which are
@@
To see the connection with catastrophe theory, we need to evaluate $V_{L}$ in the vicinity of the EPs. Since $V_{L}$ only has terms that are linear in the control parameters, we can reduce the dimensions of the parameter space by rescaling by one of the parameters, which we choose to be $\gamma$. The parameter space for the EP analysis, $(\alpha/\gamma,\beta/\gamma,\delta/\gamma)$, is thus three-dimensional. In the linear limit ($\alpha/\gamma=0$), the matrix $H$ has two separate EPs (bifurcation points) in the $(\beta/\gamma,\delta/\gamma)$ plane at $(\beta/\gamma,\delta/\gamma)=(\pm 1,0)$. The coalescing eigenvector element (degenerate critical point) is $(w,\theta)=(1,3\pi/2)$ for $\beta/\gamma=1$ and $(w,\theta)=(1,\pi/2)$ for $\beta/\gamma=-1$. Since the critical point behavior around the two linear EPs is related by reflection symmetry about the plane $\beta=0$, it suffices to analyze one, $\beta/\gamma=1$, of the two. We choose, for convenience, the $(w,\theta)$-independent terms in Eq. (4) for $V_{L}$ to shift $V_{L}$ to zero at the $\beta/\gamma=1$ linear EP. We firstly consider the properties of the nonlinear EP set in the immediate neighborhood of this linear EP. We write $r=w-1$, $\phi=\theta-3\pi/2$, $\beta^{\prime}/\gamma=\beta/\gamma-1$ and denote by $V_{EU}(r,\phi,\alpha/\gamma,\beta^{\prime}/\gamma,\delta/\gamma)$ the lowest-order terms in an expansion in $(r,\phi,\alpha/\gamma,\beta^{\prime}/\gamma,\delta/\gamma)$ of $V_{L}$. Then, we find
@@
[…]
@@
The form of Eq. (5) matches that of a universal unfolding of the $D_{-4}$ singularity, also called elliptic umbilic singularity, around the linear EP. Its degenerate critical points and bifurcations, which are well understood, are obtained as simultaneous solutions to three equations: $\partial V_{EU}/\partial r=0$, $\partial V_{EU}/\partial\theta=0$, and $\det\{h\}=0$, where $h$ is the Hessian matrix of $V_{EU}$. The singularity set, plotted in Fig. 2, consists of a pair of two conical surfaces with deltoid cross sections, the apices of which meet at the linear EPs at $(\alpha,\beta^{\prime},\delta)=(0,0,0)$ and $(0,-2\gamma,0)$. There are four and two critical points in state variable space at each control parameter point inside and outside the cones, respectively; also see Fig. 3 below. The ribs of each cone are cusp points where three critical points inside the cone meet and emerge as one critical point outside the cone. Each point on the surface between two ribs is a fold point where two critical points inside the cone meet and annihilate each other, thus vanishing on the outside. This arrangement of the fold and cusp singularities is characteristic of the elliptic umbilic catastrophe ($D_{-4})$. We call the meeting point of the cones the ‘organizing point’ of the catastrophe. We show in Fig. 3 that the cone-like topological structure is not limited to the immediate neighborhood of the linear EP, but remains valid even at large nonlinearities (large $\alpha$).
@@
Regarding PT symmetry hassan-etal.2015 ; miri-alu.2016 ; ge-elganainy.2016 ; kominis-etal.2017 ; teimourpour-etal.2017 ; xia-etal.2021 ; lee-etal.2024 ; klauck-etal.2025 ; bender-hook.2024 , we note that the Hamiltonian in Eq. (1) is PT-symmetric, i.e. $[H,\mathrm{PT}]=0$, with parity represented by the $\sigma_{1}$ Pauli matrix, if $\delta=0$. In that case, PT symmetry is unbroken (i.e. the eigenenergies are real and the eigenvectors are common to $H$ and PT) if $|\beta|>\sqrt{\alpha^{2}+\gamma^{2}}$. This region is indicated as thick blue line in Fig. 3.
@@
The power of using catastrophe theory in the context of EPs can be seen from the following considerations. First, the information on the geometric structure of the nonlinear EPs provides immediate knowledge of the number of eigenvalues as a function of the parameters used. The potential $V_{EU}$ has four critical points inside the cone and two outside the cone; exemplary plots of $V_{EU}$ are shown in the SM. This implies that we have four eigenvalues inside the cone and two outside, and the number changes at the folds and cusps, which represent the location of the EPs. This is illustrated in Fig. 4, which shows the eigenvalues of $H$ along the straight lines in ($\delta,\beta^{\prime})$-parameter space as indicated in each row of subplots. At the apices of the cones, Fig. 2, the four critical points coalesce, showing that, depending on how the limit is performed, the linear EP can be viewed as the limiting case of a 4th-order nonlinear EP. A stability analysis of the solutions is planned as future work.
@@
Figure 4: Complex eigenvalues (real and imaginary parts) of $H$ in Eq (1) at fixed, nonzero $\alpha/\gamma=1$ across lines in the $(\beta^{\prime},\delta)$ plane shown as insets. Regardless of the lines’ directions, we have two and four eigenvalues outside and inside the cone, respectively, in agreement with the number of values of critical points of the Lyapunov potential (see SM for examples of the potential landscape). At the smooth parts of the deltoid boundary (fold lines), two states are created/annihilated, and two states are unaffected by the boundary (a,b,e,f). Crossing the cusp, (c,d), three states coalesce at a 3rd-order nonlinear EP, denoted as EP3, and one state is unaffected.
The second aspect highlighting the power of catastrophe theory lies in the fact that the linear matrix fixes already the elliptic umbilic catastrophe, and a broad class of matrices extending our case given in Eq. 1 will have the same singularity structure. For example, we show in the SM that any diagonal linear or nonlinear perturbation that depends only on $w$ through a function $F(w)$ with $F(w)=0$ and $dF(w)/dw\neq 0$ at the linear EP is in this category. It also includes the case of higher-order nonlinearities and additional linear parameters on the diagonal matrix elements. Furthermore, relaxing the assumption of a symmetric matrix and/or adding linear parameters to the off-diagonal matrix elements is allowed. Higher-power (smooth) nonlinearities $F(w)$ may limit the validity of the exact elliptic umbilic structure to a finite region around the linear EP. For the class of $2\times 2$ matrices delineated in this paper, this local validity is always maintained.

### v3

@@ 1.1 Example: quadratic nonlinearity
where $a=\delta+i\gamma+\alpha\left(|x|^{2}-|y|^{2}\right)$, $\alpha,\beta,\gamma,\delta\in\mathbb{R}$ are parameters, and all variables and parameters are assumed scaled and unitless. This parametric form is chosen for ease of analysis and the specific form of the nonlinearity includes a broad range of physical systems, for example photonic systems with Kerr-type nonlinearity, atomic and polariton condensates, and superfluids [86, 87, 88, 89]; the restrictions can be relaxed and the method is still applicable (see Sec. 1.2 and Sec. I of the Supplementary Information for details). In the linear limit, the two states are separated by $2\delta$ in frequency, and one state gains and the other dissipates at equal rates $|\gamma|$.
Eliminating the eigenvalue $E$ from Eq. (2) reduces the matrix equation to a nonlinear equation in the complex eigenvector element $\tilde{w}$, the real and imaginary parts of which are
@@
To see the connection with catastrophe theory, we need to consider $V_{L}$ in the vicinity of the linear EPs. Since $V_{L}$ only has terms that are linear in the control parameters, we can reduce the dimensions of the parameter space by rescaling by one of the parameters, which we choose to be $\gamma$. The parameter space for the EP analysis, $(\alpha/\gamma,\beta/\gamma,\delta/\gamma)$, is thus three-dimensional. In the linear limit ($\alpha/\gamma=0$), the matrix $H$ has two separate EPs (bifurcation points) in the $(\beta/\gamma,\delta/\gamma)$ plane at $(\beta/\gamma,\delta/\gamma)=(\pm 1,0)$. The coalescing eigenvector element (degenerate critical point) is $(w,\theta)=(1,3\pi/2)$ for $\beta/\gamma=1$ and $(w,\theta)=(1,\pi/2)$ for $\beta/\gamma=-1$. Since the critical point behavior around the two linear EPs is related by reflection symmetry about the plane $\beta=0$, it suffices to analyze one, $\beta/\gamma=1$, of the two. We choose, for convenience, the $(w,\theta)$-independent terms in Eq. (5) for $V_{L}$ to shift $V_{L}$ to zero at the $\beta/\gamma=1$ linear EP. We firstly consider the properties of the nonlinear EP set in the immediate neighborhood of this linear EP. We write r=w1r=w-1, $\phi=\theta-3\pi/2$, $\beta^{\prime}/\gamma=\beta/\gamma-1$ and denote by $V_{EU}(r,\phi,\alpha/\gamma,\beta^{\prime}/\gamma,\delta/\gamma)$ the lowest-order terms in an expansion in $(r,\phi,\alpha/\gamma,\beta^{\prime}/\gamma,\delta/\gamma)$ of $V_{L}$. Then, we find
@@
$={}$ $-\left[\frac{\phi^{3}}{3}-\phi r^{2}+\frac{\alpha}{\gamma}r^{2}-2\frac{\beta^{\prime}}{\gamma}\phi-2\frac{\delta}{\gamma}r\right]\,.$
The form of Eq. (6) matches that of a universal unfolding of the $D_{-4}$ singularity, also called elliptic umbilic singularity, around the linear EP. Its degenerate critical points and bifurcations, which are well understood, are obtained as simultaneous solutions to three equations: $\partial V_{EU}/\partial r=0$, $\partial V_{EU}/\partial\theta=0$, and $\det\{h\}=0$, where $h$ is the Hessian matrix of $V_{EU}$. The corresponding set of three equations, with $V_{L}$ taking the place of $V_{EU}$, give the EP set in the whole parameter space (the singularity set). Since there are five unknowns, $w$, $\theta$ , $\alpha/\gamma$, $\beta^{\prime}/\gamma$, $\delta/\gamma$, the solution set of the three equations forms a 2D surface in the 5D combined state variable and parameter space. The projection of this solution surface onto the parameter space is the singularity set. Algebraic detail of the solution set can be found in Sec. III of the Supplementary Information .
@@
The singularity set, plotted in Fig. 2, consists of a pair of two conical surfaces with deltoid cross sections, the apices of which meet at the linear EPs at $(\alpha,\beta^{\prime},\delta)=(0,0,0)$ and $(0,-2\gamma,0)$. There are four and two critical points in state variable space at each control parameter point inside and outside the cones, respectively; also see Fig. 3 below. The ribs of each cone are cusp points where three critical points inside the cone meet and emerge as one critical point outside the cone. Each point on the surface between two ribs is a fold point where two critical points inside the cone meet and annihilate each other, thus vanishing on the outside. This arrangement of the fold and cusp singularities is characteristic of the elliptic umbilic catastrophe ($D_{-4})$. We call the meeting point of the cones the ‘organizing point’ of the catastrophe. We show in Fig. 3 that the cone-like topological structure is not limited to the immediate neighborhood of the linear EP, but remains valid even at large nonlinearities (large $\alpha$). More information on the EP set and the Lyapunov potential is given in Sec. IV and V of the Supplementary Information.
Interestingly, we see that the non-trivial topology of the set of EPs in the nonlinear parameters space (the parameter space that includes the nonlinearity $\alpha$, as shown in Fig. 2), is already fully determined by the EPs in the linear space (the subspace of parameters where the nonlinearity is zero). In other words, there are infinitely many possibilities for the EP topology in the nonlinear case (cf. Fig. 1b), but already from the properties of the linear EP alone we find that the topology must be the cone-like structure. At small nonlinearities, this structure is of the form of an elliptic umbilic, i.e. a cone with a deltoid cross section (and the cone-like structure can persist at larger nonlinearities, as in our example, preserving the topology of the elliptic umbilic but not its precise geometry). This is a remarkable statement, since usually one would try to find the set of EPs in the nonlinear parameters space by searching or scanning the part of the parameter space where the nonlinearity is non-zero. To be more specific, if in Eq. (6) we set $\alpha=0$, $\beta^{\prime}=0$ and $\delta=0$, we are left with the so-called ‘germ’ (sometimes called degenerate core) of the catastrophe. For catastrophes with codimension up to 5 (called elementary catastrophes), the germs can be found in textbooks, e.g. [33, 75] and are reproduced in Supplementary Table I. Each germ gives an unambiguous assignment of the corresponding elementary catastrophe. Since, in our case, the Lyapunov potential (as a polynomial in the state variables given by the eigenvectors) that generates the eigenvector equations of the nonlinear eigenvalue problem exists, we can predict the topology of the EP singularity set in the full parameter space solely from the Lyapunov potential where all parameters (including the nonlinearity) have been set to zero. The precise geometric shape can change under smooth transformations of parameters, but its topological properties are determined by the universal unfolding (which is defined as the versal unfolding of minimal dimension, see, e.g., p. 621 of Ref. [75]) of the elliptic umbilic singularity.
@@
Regarding PT symmetry [36, 37, 38, 39, 40, 41, 42, 43, 44, 35], we note that the Hamiltonian in Eq. (2) is PT-symmetric, i.e. $H,\mathrm{PT}]=0$, with parity represented by the $\sigma_{1}$ Pauli matrix, if $\delta=0$. In that case, PT symmetry is unbroken (i.e. the eigenenergies are real and the eigenvectors are common to $H$ and PT) if $|\beta|>\sqrt{\alpha^{2}+\gamma^{2}}$. This region is indicated as thick blue line in Fig. 3. In Fig. 2, it is the part of the $\delta=0$ plane outside of the deltoid cone.
@@
The power of using catastrophe theory in the context of EPs can be seen from the following considerations. First, the information on the geometric structure of the nonlinear EPs provides immediate knowledge of the number of eigenvalues as a function of the parameters used. The potential $V_{EU}$ has four critical points inside the cone and two outside the cone; exemplary plots of $V_{EU}$ are shown in Supplementary Figure S3 the Supplementary Information. This implies that we have four eigenvalues inside the cone and two outside, and the number changes at the folds and cusps, which represent the location of the EPs. This is illustrated in Fig. 4, which shows the eigenvalues of $H$ along the straight lines in ($\delta,\beta^{\prime})$-parameter space as indicated in each row of subplots. At the apices of the cones, Fig. 2, the four critical points coalesce, showing that, depending on how the limit is performed, the linear EP can be viewed as the limiting case of a 4th-order nonlinear EP. A stability analysis of the solutions is planned as future work.
@@
Figure 4: Selected complex eigenvalue traces. Complex eigenvalues (real and imaginary parts) of $H$ in Eq (2) at fixed, nonzero $\alpha/\gamma=1$ across lines in the $(\beta^{\prime},\delta)$ plane shown as insets. Regardless of the lines’ directions, we have two and four eigenvalues outside and inside the cone, respectively, in agreement with the number of values of critical points of the Lyapunov potential (see Supplementary Figure S3 for examples of the potential landscape). At the smooth parts of the deltoid boundary (fold lines), two states are created/annihilated, and two states are unaffected by the boundary (a,b,e,f). Crossing the cusp, (c,d), three states coalesce at a 3rd-order nonlinear EP, denoted as EP3, and one state is unaffected.
1.2 Extension beyond quadratic nonlinearities
In the previous section, Sec. 1.1, we discussed the example of a Kerr-like nonlinearity. We note, however, that the power of catastrophe theory lies in the fact that the linear matrix already fixes the elliptic umbilic catastrophe, and a broad class of matrices extending our case given in Eq. (2) will have the same topological singularity structure. Let us first note that the matrix in Eq. (2) is equivalent, as far as the topology of the EP object in parameter space is concerned, to one that contains self coupling ($\alpha_{c}$) and cross coupling ($\alpha_{x}$) terms with quadratic nonlinearities:
$H_{1}=\begin{pmatrix}\delta+i\gamma+\alpha_{c}|x|^{2}+\alpha_{x}|y|^{2}&\beta\\ \beta&-\delta-i\gamma+\alpha_{c}|y|^{2}+\alpha_{x}|x|^{2}\end{pmatrix}\,$. (7)
Here, the parameter $\alpha$ used above corresponds to $\alpha=(\alpha_{c}-\alpha_{x})/2$. The two matrices differ by a multiple of the unit matrix, and thus have the same eigenvectors while the eigenvalues differ only by a constant shift $\bar{\alpha}=(\alpha_{c}+\alpha_{x})/2$. Here, we assume the couplings $\alpha_{c}$ and $\alpha_{x}$ to be real-valued, but below we will see that they also can be imaginary (which in some applications can be related to gain saturation) or complex. We can generalize the matrix in Eq. (7) to allow for sufficiently smooth real valued nonlinearity functions $f(u)$ (for which conditions will be given below) and with complex nonlinearity parameters $\alpha_{c}$, $\alpha_{x}$
$H_{1}=\begin{pmatrix}\delta+i\gamma+\alpha_{c}f(|x|)+\alpha_{x}f(|y|)&\beta\\ \beta&-\delta-i\gamma-\alpha_{x}f(|x|)+\alpha_{c}f(|y|)\end{pmatrix}\,$. (8)
Similar to the case discussed above, we can subtract a diagonal matrix that is a unit matrix multiplied by the (nonlinear) constant $g(\mathrm{\bf x})=\bar{\alpha}(f(|x|)+f(|y|))$, i.e. we define $H_{1}=g(\mathrm{\bf x})\hat{1}+H$, where $\hat{1}$ is the unit matrix, and obtain
$H=\begin{pmatrix}\delta+i\gamma+\alpha(f(|x|)-f(|y|)&\beta\\ \beta&-\delta-i\gamma-\alpha(f(|x|)-f(|y|))\end{pmatrix}\,$. (9)
Once the nonlinear eigenvalue problem of $H$ has been solved and the eigenvalues $\lambda_{n}$ and eigenvectors $\mathrm{\bf x}_{n}$ are known, we evaluate the constant $g(\mathrm{\bf x})$ with the eigenvectors $\mathrm{\bf x}_{n}$ and obtain the eigenvalues of $H_{1}$ in Eq. (8) as $\lambda_{n}^{(1)}$ = $\lambda_{n}+g(\mathrm{\bf x}_{n})$, while the eigenvectors for $H_{1}$ and $H$ are the same. As in the previous section, we write the eigenvector elements as functions of $w$ and $\theta$, and note that the moduli of both components, $|x|=1/\sqrt{1+w^{2}}$ and $|y|=w/\sqrt{1+w^{2}}$ , only depend on the real variable $w$. Hence the nonlinear term can be written as
$\alpha(f(|x|)-f(|y|)=\alpha F(w)$ (10)
with
$F(w)=f(\frac{1}{\sqrt{1+w^{2}}})-f(\frac{w}{\sqrt{1+w^{2}}})$ (11)
Similar to the example of the quadratic nonlinearity discussed in the previous section, the EP is at $w=1$. As mentioned at the beginning of Sec. 1, the condition for the elliptic umbilic singularity is that at the linear EP (where $w=1$ in the case of symmetric matrices), the function $F(w)$ must be approximated by a power expansion. Therefore, important examples of functions with elliptic umbilic singularity sets include all power functions $f(u)=u^{n}$ with nonzero exponent, $n\neq 0$, such as the quartic nonlinearity $f(u)=u^{4}$, and periodic functions such as $f(u)=\cos(u^{2})$.
1.3 Different Taylor expansions
So far, we have limited the analysis to symmetric matrices where, at the EP (i.e. for $w=1)$ the Taylor expansion of the nonlinearity function $F(w)$ yields $F(1)=0$ and $F^{\prime}(1)\neq 0$. We note, however, that the theory remains valid for any real-valued function $F(w)$ that can be approximated by a Taylor expansion in the vicinity of the EP. In this context, we consider separately four categories of functions that yield elliptic umbilic EP structures that include and extend the example discussed in Sec. 1.1 and 1.2. These are (1) $F(1)\neq 0$ and $F^{\prime}(1)\neq 0$, (2) $F(1)=0$ and $F^{\prime}(1)\neq 0$, (3) $F(1)\neq 0$ and $F^{\prime}(1)=0$, and (4) $F(1)=0$ and $F^{\prime}(1)=0$. Here category (2) is the one discussed above. We note that in the four categories the elliptic umbilic singularity sets emerge as functions of different parameters/variables. But common to these categories is the elliptic umbilic topology, in other words their singularity sets are topologically equivalent. Further details about the Lyapunov functions for the four categories are given in Sec I.B.1 of the Supplementary Information. Examples of matrices belonging to category (1) include the Hamiltonian in Eq. (1) of Ref. [66] evaluated without noise, and Eq. (5) of Ref. [34].
1.4 Asymmetric matrices, complex nonlinear coupling constants
We have also considered the case of asymmetric matrices, with the off-diagonal parameters $b$ and $c$ in the Hamiltonian Eq. (1) being independent complex parameters. Our findings are summarized here, and the algebraic derivation is detailed in Sec. I.B.3 and I.B.4 in the Supplementary Information. As in the symmetric-matrix case, we construct a Lyapunov potential for the eigenvector elements of this asymmetric Hamiltonian in Sec. I.B.3. A straightforward expansion around the linear EP shows that the germ (the limiting Lyapunov potential as a function of ($r$, $\phi$) when the parameters are set at the EP values) has the structure of a double elliptic umbilic. We derive in Sec. I.B.4 a coordinate transformation that maps this structure into a canonical elliptic umbilic in the new coordinates. This shows that the (linear and nonlinear) EP sets also possess the topological structure of the elliptic umbilic around the EP.
We find that the elliptic umbilic topology is also retained when the nonlinear coupling constant $\alpha$ is imaginary or complex. We comment briefly on this case in Sec. I.B.2 in the Supplementary Information.
1.5 Symmetry in nonlinear coefficient $\alpha$
We also mention again that the EP structure shown in Fig. 2 is symmetric with respect to the Kerr-like nonlinearity $\alpha$ and the linear coupling $\beta$ (i.e., invariant under $\alpha\rightarrow-\alpha$, $\beta\rightarrow-\beta$). For real-valued self ($\alpha_{c}$) and cross ($\alpha_{x}$) couplings in Eq. (7), it is instructive to compare our nonlinear $2\times 2$ eigenvalue problem with the nonlinear Schrödinger equation (NLS). If the NLS is derived from the Helmholtz equation in paraxial approximation, then the spatial propagation coordinate $z$ plays the role of time and the diffraction (second-order derivative in the transverse direction) plays the role of the kinetic energy, see e.g. [90]. While the general solution structure of the NLS includes effects like soliton formation and is generally complicated (and the NSL is not integrable in more than one spatial dimensions), it has been shown in [91] that coupled-mode analysis in the nonlinear regime can be done with the NLS under suitable conditions. If one would use two modes with different transverse extensions, effects analogous to self-focusing and self-defocusing could be described by a 2-mode model. In the language of our eigenvalue equation, the eigenvalue $E$ would correspond to the wave vector in the propagation direction $z$, and the coupling coefficients $\alpha_{c}$ and $\alpha_{x}$ would depend on the spatial overlap between the two modes. More detailed studies of the discrete NLS (also called discrete self-trapping equation) are provided in Ref. [92], but most studies of the discrete NLS, including 2-mode models that are equivalent to the (nonlinear) pendulum [93], do not include decay, which is essential in our study of the EP structure. If, however, a self-focusing and self-defocusing system with decay can be described by our 2-mode model, Eq. (7), with the difference in the absorption coefficients of the two modes given by $\gamma$, then we predict that the EP structure is symmetric in $\alpha=\frac{1}{2}(\alpha_{c}-\alpha_{x})$ as well as the linear coupling $\beta$ and the energy splitting $\delta$, as shown in Fig. 2.
1.6 Larger matrix dimensions